1. Two charges are 60cm apart, in air. One charge Q1 is +1.67 x 10-7C; The other Q2 is

-1.67 x 10-7 C. What is the electric field intensity at P midway between the charges?


2. The average distance between the electron and the proton in a certain atom is 5 x 10-10m. How strong is the electric field that the proton experiences?


3. Two charges Q1 = 2.3 x 10-9C and Q2 = 3.5 x 10-9C, are 55mm apart. What is the electric field halfway between them?


4. Two charges of 6uC are 2.5m apart. Where is the electric field along the line joining them equal to zero?


5. What is the electric field at one vertex of an equilateral triangle whose sides are 10cm long if there are charges of -30uC at the other vertexes?


Explain: Can the force caused by an electric field be greater than the force due to a gravitational field? Give an example to prove this.



1
Expert's answer
2022-06-02T10:26:10-0400

"E=\frac{4k(q_2-q_1)}{r^2}=14.3~\frac{kN}C."